If 15mL of a 6.5M HNO3 solution is added to 55mL of water, what is the new concentration?


1
Expert's answer
2021-02-12T05:03:19-0500
"C=n\/V"

"n(HNO3)=6.5*0.015=0.0975 mol"

"\Sigma V=15+55=70 mL=0.07 L"

"C(new)=0.0975\/0.07=1.4 M"